Can someone please help me!!!
A sample of radioactive iodine has a half-life of 27 days. (a) Determine an exponential decay model for the amount of radioactive material left in the sample after t days; and (B) Determine how long it will take for 90% of the sample to decay.
Thanks,

Answer by stanbon!(97)

Let t represent # of days during which decay occurs.
Then t/27 will tell us how many times the iodine has
decreased by 1/2.
1st: So, A = a (1/2)^(t/27), where "a" is the original amount
and A is the amount remaining after t-days.
2nd: If 90% is gone there is only 10% of "a" remaining.
So, 10%(a) = a (1/2)^(t/27)
0.1 = 2^(-t/27)
Take the log of both sides to get:
log 0.1 = (-t/27)log27
Then -t/27 = (log 0.1) / (log 27)
-t = 27(-0.699...)
t = 18.86 days
Cheers,
Stan H.